# Multi-scale Restoration of Missing Data in Optical Time-series Images with Masked Spatial-Temporal Attention Network

### v1

@@ Abstract
Due to factors such as thick cloud cover and sensor limitations, remote sensing images often suffer from significant missing data, resulting in incomplete time-series information. Existing methods for imputing missing values in remote sensing images do not fully exploit spatio-temporal auxiliary information, leading to limited accuracy in restoration. Therefore, this paper proposes a novel deep learning-based approach called MS2TAN (Multi-scale Masked Spatial-Temporal Attention Network), for reconstructing time-series remote sensing images. Firstly, we introduce an efficient spatio-temporal feature extractor based on Masked Spatial-Temporal Attention (MSTA), to obtain high-quality representations of the spatio-temporal neighborhood features in the missing regions. Secondly, a Multi-scale Restoration Network consisting of the MSTA-based Feature Extractors, is employed to progressively refine the missing values by exploring spatio-temporal neighborhood features at different scales. Thirdly, we propose a “Pixel-Structure-Perception” Multi-Objective Joint Optimization method to enhance the visual effects of the reconstruction results from multiple perspectives and preserve more texture structures. Furthermore, the proposed method reconstructs missing values in all input temporal phases in parallel (i.e., Multi-In Multi-Out), achieving higher processing efficiency. Finally, experimental evaluations on two typical missing data restoration tasks across multiple research areas demonstrate that the proposed method outperforms state-of-the-art methods with an improvement of 0.40dB/1.17dB in mean peak signal-to-noise ratio (mPSNR) and 3.77‰/9.41‰ in mean structural similarity (mSSIM), while exhibiting stronger texture and structural consistency.
@@ I Introduction
Researchers have proposed various methods to recover missing data in remote sensing images. Early approaches to missing value restoration can be broadly categorized into three types [1]: spatial-based, temporal-based, and spatio-temporal-based data recovery. These methods have shown promising results in specific scenarios with low resolution and low missing rates. However, most of them rely on linear models and struggle to handle complex and detailed scenes. Moreover, due to limited reference information, the generated images often exhibit blurriness and lack continuous textures.
@@
In recent years, the rapid advancement of deep learning theory and computer hardware [2] has led to significant progress in remote sensing image restoration using deep learning-based methods. These methods can be broadly categorized into two types: CNN-based [3] and Transformer-based [4] approaches. Compared to traditional statistical models, CNNs exhibit strong non-linear expressive power, allowing for efficient extraction of spatial features from remote sensing data and significantly improving the accuracy of image restoration. However, these structures struggle to fully exploit time-series information, resulting in a bottleneck in restoration accuracy. Transformers, based on self-attention mechanisms, possess a global receptive field, enabling comprehensive and efficient utilization of both temporal sequences and images [5] to enhance reconstruction results [6]. However, due to the high resolution and long time-series of remote sensing images, token sequences become excessively long, leading to extremely high complexity in attention computations.
@@
To efficiently mine spatio-temporal information in remote sensing images, we apply self-attention mechanism separately in the temporal and spatial dimensions and alternate between the two [7], greatly reducing computational complexity. To address the problem of significant distributional differences between missing and non-missing values in remote sensing images, we apply missing values mask and diagonal mask to the attention matrix [8], proposing Masked Spatial-Temporal Attention (MSTA) to enhance the expressive power of spatio-temporal attention and optimize the spectral discrepancy at the transition regions.
@@
For full exploitation of the spatio-temporal neighborhood features at different scales [9], we further propose a Multi-scale Restoration Network. The network consists of MSTA-based Feature Extractors with different embedding scales, which progressively refine the reconstruction of missing information from coarse to fine granularity levels, achieving higher restoration accuracy [10]. For model training, we propose a “Pixel-Structure-Perception” Multi-Objective Joint Optimization method, using pixel-wise loss as the basic loss for the restoration task and incorporating structural loss [11] and perceptual loss [12] to optimize the model’s results from the perspectives of structure, texture, shape, and spatial relations, thereby achieving high-quality image inpainting.
Finally, we conducted simulated and real data experiments to validate the performance of the proposed method for the ETM+ SLC-off and thick cloud removal tasks. The results from six research areas indicate that our method outperforms competitors. In addition, a validation study demonstrates the effectiveness of the core components.
@@
    We propose a deep learning-based approach called MS2TAN, for repairing missing information in time-series remote sensing images. Our method utilizes a Multi-scale Restoration Network with MSTA-based Feature Extractors to learn the non-linear mapping between missing and complete data. Both simulated and real data experiments show that proposed MS2TAN outperforms state-of-the-art methods across multiple research areas.
@@
    To address the challenges of high resolution and long temporal sequences in remote sensing data, we introduce the Masked Spatial-Temporal Attention (MSTA) mechanism. The MSTA efficiently extracts spatio-temporal features, improves the utilization of spatio-temporal contextual information, reduces color transition artifacts at the boundaries of missing value regions, and significantly reduces the time complexity of self-attention.
@@
The rest of this paper is organized as follows. In Section II, we review the existing methods for reconstructing missing information in remote sensing images. The network architecture and methodology details of our proposed model are presented in Section III. In Section IV, we showcase the results of missing data reconstruction in both simulated and real data experiments, compare them with mainstream methods, and validate the effectiveness of each component through validation study. Finally, our conclusions and future prospects are summarized in Section V.
@@ II-A Traditional methods
Early research on the restoration of missing data in remote sensing images can be roughly divided into three categories: spatial-based, temporal-based, and spatio-temporal hybrid methods.
@@ II-A2 Temporal-based Methods
II-A3 Spatio-temporal-based Methods
To overcome these limitations, spatio-temporal-based methods integrate the spatial and temporal correlations to reconstruct missing data under various conditions. For instance, improved nearest neighbor pixel interpolation methods [27], methods based on spatio-temporal Markov random field model [28], spatio-temporal weighted regression model [29], methods based on group sparse representation [30], and methods based on low-rank tensor decomposition [31, 32] have been proposed. However, most of these methods rely on linear models and struggle to handle complex and intricate scenes.
@@ II-B1 CNN-based methods
CNNs have shown high efficiency in extracting spatial features from remote sensing data, leading to significant improvements in the accuracy of remote sensing image restoration. Malek et al. [33] applied a context encoder [34] to reconstruct thick clouds in remote sensing images. CNNs combined with GAN structures [35, 36] were used to generate realistic reconstructed images. Zhang et al. [37] proposed a CNN-based Spatio-Temporal Spectra (STS-CNN) framework, which was further developed into a progressive spatio-temporal patch grouping framework [38]. CNNs that incorporate temporal inputs through channel-wise concatenation [39] merge feature maps of target images and temporal images, introducing auxiliary information for missing data restoration. However, CNNs lack a true understanding of time series, which hinders the efficient integration of temporal and spatial information in these methods and limits the utilization of long temporal sequences as auxiliary information, ultimately affecting the restoration accuracy.
II-B2 Transformer-based methods
Transformer [4] employs self-attention mechanism to process text sequences and capture global contextual information. It has been proven that Transformer-based methods, such as Visual Transformer (ViT) [5], also exhibit powerful performance in the field of image processing. The sequence modeling capability of self-attention mechanism can be leveraged to assist in high-precision restoration of missing values in remote sensing images. Xu et al. [40] applied spatial self-attention to feature maps to capture the distribution of cloud thickness. Christopoulos et al. [41] utilized axial attention to remove thick clouds in remote sensing images. Recently, Liu et al. [42] used spatial attention and channel attention to remove cloud cover in the images. Li et al. [43] introduced SAR auxiliary data and synthesized cloud-free images using a Transformer-based GAN.
However, the spatial attention used in ViT requires computing the correlations between all pairs of patches. In the context of processing time-series remote sensing images, the images often have high resolutions and long time-series, resulting in a large number of tokens and extremely high computational complexity. Bertasius et al. [7] conducted a detailed comparison of various forms of spatial-temporal attention and proposed separated spatial-temporal attention that achieves a balance between efficiency and performance. We further improved the separated spatial-temporal attention by introducing missing value masks and diagonal masks [8], resulting in Masked Spatial-Temporal Attention (MSTA), which exhibits superior performance in the task of missing value restoration. Compared to CNN-based methods and original ViT methods, MSTA enables efficient processing of spatio-temporal information and leverages long-term temporal information to assist in missing value reconstruction, leading to more precise reconstruction results.
@@ III-A1 Problem Definition
The original time-series remote sensing image with missing values can be represented as $X\in\mathbb{R}^{T\times C\times H\times W}$, where $T$ denotes the length of the time series, $C$ represents the number of channels, and $H$ and $W$ denote the spatial dimensions of the region. $X_{(t,c,i,j)}$ denotes the value of channel $c$ at position $(i,j)$ at time $t$. To account for the missing values, we introduce the missing value hint tensor $M\in\mathbb{R}^{T\times C\times H\times W}$, which is defined as shown in Equation (1). In practical applications, the missing value hint tensor $M$ can be obtained from QA band or cloud detection algorithms such as Fmask [44].
@@
The target sequence $Y$ in $\mathbb{R}^{T\times C\times H\times W}$ represents the actual data, while $\widetilde{Y}$ in $\mathbb{R}^{T\times C\times H\times W}$ represents the data repaired by the model. Therefore, the definition of the multi-temporal remote sensing image reconstruction model is given by equation (2). Here, the output $\widetilde{Y}$ of the model is the initial reconstruction result, and $\odot$ denotes the Hadamard product. By replacing the observed real values in $X$ with $\widetilde{Y}$, we obtain the final reconstruction result $\widetilde{Y}_{\mathrm{out}}$ as shown in equation (3).
@@ III-B MSTA-based Feature Extractor
The structure of the MSTA-based Feature Extractor (MFE) module is illustrated in Fig. 2. It consists of position encoding and a cascade of $L$ Masked Spatial-Temporal Attention (MSTA) units. The input to this module is a high-dimensional feature vector $E\in\mathbb{R}^{TN\times d_{\mathrm{emb}}}$, which is obtained by embedding the time-series remote sensing image $X$ and the missing information mask $M$ into blocks (as described in detail in Section III-C1). Here, $T$ represents the time-series length, $N$ represents the number of patches in a single remote sensing image, and $d_{\mathrm{emb}}$ represents the dimension of the token vector corresponding to each patch. Subsequently, the position encoding introduces spatial-temporal positional semantics, followed by $L$ layers of MSTA units, and ultimately produces a sequence with incorporated spatio-temporal correlation features.
@@ III-B1 Masked Self-Attention
The Transformer model utilizes self-attention mechanism to model sequences. It maps the input to query vector Q, key vector K, and value vector V. The attention score between Q and K is computed using Scaled Dot-Product, followed by the application of the Softmax function to obtain attention weights A. The final output is the attention-weighted V, as shown in Equation 4.
@@
  • The diagonal mask sets the diagonal of the attention matrix as $\infty$- ∞, preventing each step from attending to itself and forcing it to rely on the other $TN-1$ steps for inference. This helps capture the spatio-temporal feature correlation in high-dimensional space.
More specifically, for the attention score matrix $\mathrm{Sc}\in\mathbb{R}^{TN\times TN}$, the attention mask operation $\mathrm{ApplyMask}$ is defined as shown in Equation (5).
@@
Here, $Mask()$ checks whether the mask condition is satisfied, $\mathrm{MR}(e)$ represents the missing rate of the corresponding patch $e$ (which can be calculated from the hint tensor $M$), and $C_{\mathrm{max}}$ is a hyperparameter that controls the maximum allowable missing rate. Applying Equation (5) to the attention scores in Equation (4), we obtain the expression for a single masked attention head, as shown in Equation (6). Finally, by linearly projecting and combining multiple attention heads with the residual connection, we obtain the output of the masked self-attention, as shown in Equation (7).
$\mathrm{H}^{(i)}=\mathrm{A}^{\prime}\mathrm{V}=\mathrm{Softmax}\left(\mathrm{ApplyMask}\left(\frac{\mathrm{Q}\cdot\mathrm{K}^{\top}}{\sqrt{d_{k}}}\right)\right)\cdot\mathrm{V}$ (6)
@@
In Equation (7), $\mathrm{A}^{\prime}$ denotes the attention weights post mask application, $\mathrm{MaskedAttn}()$ represents the masked self-attention operation, $\mathrm{Proj}$ stands for projection head, and $h$ denotes the number of attention heads.
Figure 3: Schematic of MSTA. Taking the blue patch as an example, patches at the same spatial position are considered as temporal auxiliaries, while patches at the same time are regarded as spatial auxiliaries. Additionally, patches with excessively high missing rates will be excluded from the receptive field to avoid the influence of distributional differences on model learning.
@@ III-B2 Masked Spatial-Temporal Attention
Masked Spatial-Temporal Attention (MSTA) captures the spatio-temporal correlations within the input sequence, with its masking mechanism and receptive field depicted in Fig. 3. Specifically, MSTA comprises Masked Temporal Attention (MTA), Masked Spatial Attention (MSA), Layer Normalization (LN), and Feed-Forward Network (FFN). These components are elaborated upon below.
For an input sequence $E\in\mathbb{R}^{TN\times d_{\mathrm{emb}}}$, MTA first reshapes $E$ to place the time dimension and the feature dimension at the end, obtaining a sequence $e_{t}$ of length $T$ and $d_{\mathrm{emb}}$ dimensions. Subsequently, for each position in space, MaskedAttn operation is performed along the temporal direction to obtain the output of MTA, as shown in Equation (8).
@@
Similarly, after the corresponding dimension transformation of the input sequence, MaskedAttn operation is conducted along the spatial direction for each temporal image, yielding the output of MSA, as shown in Equation (9).
@@
FFN consists of two linear layers separated by a ReLU activation function, with a residual connection established between the input and output. The expression for FFN is presented in Equation (10).
@@
MSTA sequentially applies MTA and MSA to the input token sequence, exploring the spatio-temporal correlations among patches. Subsequently, FFN is used to fuse spatio-temporal features and introduce non-linear transformations, as shown in Equation (11).
@@
Figure 4: The network structure diagram of proposed MS2TAN (Multi-Scale Masked Spatial-Temporal Attention Network), which includes components such as Embedding, MFE, Unembedding, and Observed Value Replacement.

### v2

@@ Abstract
Remote sensing images often suffer from substantial data loss due to factors such as thick cloud cover and sensor limitations. Existing methods for imputing missing values in remote sensing images fail to fully exploit spatiotemporal auxiliary information, which restricts the accuracy of their reconstructions. To address this issue, this paper proposes a novel deep learning-based approach called MS2TAN (Multi-Scale Masked Spatial-Temporal Attention Network) for reconstructing time-series remote sensing images. First, we introduce an efficient spatiotemporal feature extractor based on Masked Spatial-Temporal Attention (MSTA) to capture high-quality representations of spatiotemporal neighborhood features surrounding missing regions while significantly reducing the computational complexity of the attention mechanism. Second, a Multi-Scale Restoration Network composed of MSTA-based Feature Extractors is designed to progressively refine missing values by exploring spatiotemporal neighborhood features at different scales. Third, we propose a “Pixel-Structure-Perception” Multi-Objective Joint Optimization method to enhance the visual quality of the reconstructed results from multiple perspectives and to preserve more texture structures. Finally, quantitative experimental results under multi-temporal inputs on two public datasets demonstrate that the proposed method outperforms competitive approaches, achieving a 9.76%/9.30% reduction in Mean Absolute Error (MAE) and a 0.56 dB/0.62 dB increase in Peak Signal-to-Noise Ratio (PSNR), along with stronger texture and structural consistency. Ablation experiments further validate the contribution of the core innovations to imputation accuracy.
@@ I Introduction
Researchers have proposed various methods to recover missing data in remote sensing images. Early approaches to missing value restoration can be broadly categorized into three types [1]: spatial-based, temporal-based, and spatiotemporal-based data recovery. These methods have shown promising results in specific scenarios with low resolution and low missing rates. However, most of them rely on linear models and struggle to handle complex and detailed scenes. Moreover, due to limited reference information, the generated images often exhibit blurriness and lack continuous textures.
@@
In recent years, the rapid advancement of deep learning theory and computer hardware [2] has led to significant progress in remote sensing image restoration using deep learning-based methods. These methods can be broadly categorized into two types: CNN-based [3] and ViT-based [4] approaches. Compared to traditional statistical models, CNNs exhibit strong non-linear expressive power, allowing for efficient extraction of spatial features from remote sensing data and significantly improving the accuracy of image restoration. However, these structures struggle to fully exploit time-series information, resulting in a bottleneck in restoration accuracy. ViTs, based on self-attention mechanism [5], possess a global receptive field, enabling comprehensive and efficient utilization of both temporal sequences and images to enhance reconstruction results [6]. However, due to the high resolution and long time-series of remote sensing images, token sequences become excessively long, leading to extremely high complexity in attention computations.
@@
To efficiently mine spatiotemporal information in remote sensing images, we apply self-attention mechanism separately in the temporal and spatial dimensions and alternate between the two [7], greatly reducing computational complexity. To address the problem of significant distributional differences between missing and non-missing values in remote sensing images, we apply missing values mask and diagonal mask to the attention matrix [8], proposing Masked Spatial-Temporal Attention (MSTA) to enhance the expressive power of spatiotemporal attention and optimize the spectral discrepancy at the transition regions.
@@
For full exploitation of the spatiotemporal neighborhood features at different scales [9], we further propose a Multi-scale Restoration Network. The network consists of MSTA-based Feature Extractors with different embedding scales, which progressively refine the reconstruction of missing information from coarse to fine granularity levels, achieving higher restoration accuracy [10]. For model training, we propose a “Pixel-Structure-Perception” Multi-Objective Joint Optimization method, using pixel-wise loss as the basic loss for the restoration task and incorporating structural loss [11] and perceptual loss [12] to optimize the model’s results from the perspectives of structure, texture, shape, and spatial relations, thereby achieving high-quality image inpainting.
Finally, we performed both quantitative and qualitative experiments on two datasets, comparing our approach with several mainstream methods. Additionally, we carried out ablation studies on the key innovations and analyzed the trade-offs between effectiveness and efficiency across models of different sizes.
@@
    We propose a deep learning-based method MS2TAN for reconstructing missing data in time series remote sensing images. Our method utilizes a multi-scale restoration network to learn an end-to-end mapping between incomplete and complete image sequences. MS2TAN achieved higher restoration accuracy than mainstream methods in quantitative experiments, and showed better visual effects in eliminating gaps in real data.
@@
    To address the challenges of high resolution and long time series in remote sensing data, we introduce the Masked Spatial-Temporal Attention (MSTA) mechanism. MSTA effectively extracts spatiotemporal features, improves the utilization of spatiotemporal context information, reduces color transition artifacts at the boundaries of missing value areas, and significantly reduces the computational complexity of self-attention.
@@
The rest of this paper is organized as follows. In Section II, we review the existing methods for reconstructing missing information in remote sensing images. The network architecture and methodology details of our proposed model are presented in Section III. In Section IV, we showcase the results of missing data reconstruction in both quantitative and qualitative experiments, compare them with mainstream methods, and validate the effectiveness of each component through validation studies. Finally, our conclusions and future prospects are summarized in Section V.
@@ II-A Traditional methods
Early research on the restoration of missing data in remote sensing images can be roughly divided into three categories: spatial-based, temporal-based, and spatiotemporal hybrid methods.
@@ II-A2 Temporal-based Methods
II-A3 Spatiotemporal Hybrid Methods
To overcome these limitations, spatiotemporal hybrid methods integrate the spatial and temporal correlations to reconstruct missing data under various conditions. For instance, improved nearest neighbor pixel interpolation methods [27], methods based on spatiotemporal Markov random field model [28], spatiotemporal weighted regression model [29], methods based on group sparse representation [30], and methods based on low-rank tensor decomposition [31, 32] have been proposed. However, most of these methods rely on linear models and struggle to handle complex and intricate scenes.
@@ II-B1 CNN-based methods
CNNs have shown high efficiency in extracting spatial features from remote sensing data, leading to significant improvements in the accuracy of remote sensing image restoration. Malek et al. [33] applied a context encoder [34] to reconstruct thick clouds in remote sensing images. CNNs combined with GAN structures [35, 36] were used to generate realistic reconstructed images. Zhang et al. [37] proposed a CNN-based spatiotemporal Spectra (STS-CNN) framework, which was further developed into a progressive spatiotemporal patch grouping framework [38]. CNNs that incorporate temporal inputs through channel-wise concatenation [39] merge feature maps of target images and temporal images, introducing auxiliary information for missing data restoration. Stucker et al. [40] used temporal self-attention for CNN’s feature map sequences in conjunction with U-Net to repair temporal images. However, CNNs lack a true understanding of time-series, which hinders the efficient integration of temporal and spatial information in these methods and limits the utilization of long temporal sequences as auxiliary information, ultimately affecting the restoration accuracy.
II-B2 ViT-based methods
Visual Transformer (ViT) [4] has shown excellent performance in many tasks in the vision domain. Xu et al. [41] applied spatial self-attention to feature maps to capture the distribution of cloud thickness. Christopoulos et al. [42] utilized axial attention to remove thick clouds in remote sensing images. Recently, Liu et al. [43] used spatial attention and channel attention to remove cloud cover in the images. However, the self-attention used in ViT requires computing the correlations between all pairs of patches (including all times). In the context of processing time-series remote sensing images, the images often have high resolutions and long time-series, resulting in a large number of tokens and extremely high computational complexity.
Bertasius et al. [7] conducted a detailed comparison of various forms of spatial-temporal attention and proposed separated spatial-temporal attention that achieves a balance between efficiency and performance. We further improved the separated spatial-temporal attention by introducing missing value masks and diagonal masks [8], resulting in Masked Spatial-Temporal Attention (MSTA), which exhibits superior performance in the task of missing value restoration. Compared to CNN-based methods and original ViT methods, MSTA enables efficient processing of spatiotemporal information and leverages long-term temporal information to assist in missing value reconstruction, leading to more precise reconstruction results.
@@ III-A1 Problem Definition
The original time-series remote sensing image with missing values can be represented as $X\in\mathbb{R}^{T\times C\times H\times W}$, where $T$ denotes the length of the time-series, $C$ represents the number of channels, and $H$ and $W$ denote the spatial dimensions of the region. $X_{(t,c,i,j)}$ denotes the value of channel $c$ at position $(i,j)$ at time $t$. To account for the missing values, we introduce the missing value hint tensor $M\in\mathbb{R}^{T\times C\times H\times W}$, which is defined as shown in Eq. (1). In practical applications, the missing value hint tensor $M$ can be obtained from QA band or cloud detection algorithms such as Fmask [44] and S2Cloudless [45].
@@
The target sequence $Y$ in $\mathbb{R}^{T\times C\times H\times W}$ represents the actual data, while $\widetilde{Y}$ in $\mathbb{R}^{T\times C\times H\times W}$ represents the data repaired by the model. Therefore, the definition of the multi-temporal remote sensing image reconstruction model is given by equation (2). Here, the output $\widetilde{Y}$ of the model is the initial reconstruction result, and $\odot$ denotes the Hadamard product. By replacing the observed real values in $X$ with $\widetilde{Y}$, we obtain the final reconstruction result $\widetilde{Y}_{\mathrm{out}}$ as shown in Eq. (3).
@@ III-B MSTA-based Feature Extractor
The structure of the MSTA-based Feature Extractor (MFE) module is illustrated in Fig. 2. It consists of position encoding and a cascade of $L$ Masked Spatial-Temporal Attention (MSTA) units. The input to this module is a high-dimensional feature vector $E\in\mathbb{R}^{TN\times d_{\mathrm{emb}}}$, which is obtained by embedding the time-series remote sensing image $X$ and the missing information mask $M$ into blocks (as described in detail in Section III-C1). Here, $T$ represents the time-series length, $N$ represents the number of patches in a single remote sensing image, and $d_{\mathrm{emb}}$ represents the dimension of the token vector corresponding to each patch. Subsequently, the position encoding introduces spatial-temporal positional semantics, followed by $L$ layers of MSTA units, and ultimately produces a sequence with incorporated spatiotemporal correlation features.
@@ III-B1 Masked Self-Attention
The Transformer model utilizes self-attention mechanism to model sequences. It maps the input to query vector Q, key vector K, and value vector V. The attention score between Q and K is computed using Scaled Dot-Product, followed by the application of the Softmax function to obtain attention weights A. The final output is the attention-weighted V, as shown in Eq. 4.
@@
  • The diagonal mask sets the diagonal of the attention matrix as $\infty$- ∞, preventing each step from attending to itself and forcing it to rely on the other $TN-1$ steps for inference. This helps capture the spatiotemporal feature correlation in high-dimensional space.
More specifically, for the attention score matrix $\mathrm{Sc}\in\mathbb{R}^{TN\times TN}$, the attention mask operation $\mathrm{ApplyMask}$ is defined as shown in Eq. (5).
@@
Here, $Mask()$ checks whether the mask condition is satisfied, $\mathrm{MR}(e)$ represents the missing rate of the corresponding patch $e$ (which can be calculated from the hint tensor $M$), and $C_{\mathrm{max}}$ is a hyperparameter that controls the maximum allowable missing rate. Applying Eq. (5) to the attention scores in Eq. (4), we obtain the expression for a single masked attention head, as shown in Eq. (6). Finally, by linearly projecting and combining multiple attention heads with the residual connection, we obtain the output of the masked self-attention, as shown in Eq. (7).
$\mathrm{H}^{(i)}=\mathrm{A}^{\prime}\mathrm{V}=\mathrm{Softmax}\left(\mathrm{ApplyMask}\left(\frac{\mathrm{Q}\cdot\mathrm{K}^{\top}}{\sqrt{d_{\text{qkv}}}}\right)\right)\cdot\mathrm{V}$ (6)
@@
In Eq. (7), $\mathrm{A}^{\prime}$ denotes the attention weights post mask application, $\mathrm{MaskedAttn}()$ represents the masked self-attention operation, $\mathrm{Proj}()$ stands for projection head, and $h$ denotes the number of attention heads.
@@ III-B2 Masked Spatial-Temporal Attention
Masked Spatial-Temporal Attention (MSTA) is used to capture the spatiotemporal correlations within the input sequence, and consists of Masked Temporal Attention (MTA), Masked Spatial Attention (MSA), Layer Normalization (LN), and Feed-Forward Network (FFN). These components are elaborated upon below.
For an input sequence $E\in\mathbb{R}^{TN\times d_{\mathrm{emb}}}$, MTA first reshapes $E$ to place the time dimension and the feature dimension at the end, obtaining a sequence $e_{t}$ of length $T$ and $d_{\mathrm{emb}}$ dimensions. Subsequently, for each position in space, MaskedAttn operation is performed along the temporal direction to obtain the output of MTA, as shown in Eq. (8).
@@
Similarly, after the corresponding dimension transformation of the input sequence, MaskedAttn operation is conducted along the spatial direction for each temporal image, yielding the output of MSA, as shown in Eq. (9).
@@
FFN consists of two linear layers separated by a ReLU activation function, with a residual connection established between the input and output. The expression for FFN is presented in Eq. (10).
@@
MSTA sequentially applies MTA and MSA to the input token sequence, exploring the spatiotemporal correlations among patches. Subsequently, FFN is used to fuse spatiotemporal features and introduce non-linear transformations, as shown in Eq. (11).
@@
Figure 3: The network structure diagram of proposed Multi-scale Restoration Network (MS2TAN), which contains several residual-connected restoration modules consisting of Embedding, MFE, and Unembedding.